Opened 4 years ago

Closed 3 years ago

Last modified 3 years ago

#585 closed flight processing (fixed)

EUFAR15/38, flight day 169b/2015, Mallorca

Reported by: lah Owned by:
Priority: immediate Milestone: 2015 data processing completion
Component: Processing: general Keywords:
Cc: Other processors:

Description

Data location: ~arsf/arsf_data/2015/flight_data/spain/EUFAR15_38-2015_169b_Mallorca

Data arrived from ARSF via SATA disk on 09/07/2015

Scientific objective: Use airborne hyperspectral, LiDAR and SfM to improve process understanding by monitoring and modelling changing patterns of ecogeomorphological connectivity in Mediterranean insular catchments.

Priority: Unknown

PI: Joan Estrany
EUFAR Project ID: MEDhy2CON

Sensors:

  • Fenix (requested)
  • Leica LIDAR (requested)
  • Leica FW LIDAR (not requested but flown anyway)
  • Owl (not requested, only available for 2 lines)

Change History (39)

comment:1 Changed 4 years ago by lah

unpacking
Have moved fenix line 1 into day 170, as date stamp was for this day, even though it is not marked on either logsheet.

comment:2 Changed 4 years ago by asm

Getting ready
Logsheet generated. There are 13 flightlines. This is in accordance with the previous entry.

Last edited 4 years ago by asm (previous) (diff)

comment:3 Changed 4 years ago by asm

Navigation Processing
Started navigation processing.

Need to use teqc to do the basestation verification.

comment:4 Changed 4 years ago by asm

Navigation Processing
Basestation verification.
(Surprisingly, exactly the same that the 169a 2015)

Lat: 39 32 26.25388
Lon: 2 43 19.99818
Elipsoidal Height: 53.556 metres
Antenna Height: 0.085 metres

comment:5 Changed 4 years ago by asm

Navigation Processing
Finished. Everything looks perfectly fine.

comment:6 Changed 4 years ago by asm

Fenix Processing
Started fenix processing.

comment:7 Changed 4 years ago by asm

Fenix Processing
Needed to set use_nav = false

Since we moved out flightline 01, there are 13 flightlines with numbers starting on 02. Meaning flightlines are named as: 02, 03, 04, .. 12, 13 ,14.

comment:8 Changed 4 years ago by asm

Hyperspectral Fenix Processing
Found SCTs values.

Flightline FENIX
1 Not on logsheet
2 2.71
3 3.54
4 3.36
5 2.73
6 2.37
7 4.08
8 12.93
9 2.59
10 2.92
11 3.60
12 3.91
13 2.62
14 20.37

Status: Awaiting fenix calibration.

Last edited 4 years ago by asm (previous) (diff)

comment:9 Changed 4 years ago by asm

Lidar Processing
Awaiting new calibration from boresight flight.

comment:10 Changed 4 years ago by asm

Lidar Processing
Started.

comment:11 Changed 4 years ago by asm

Lidar Processing

Found pitch and roll consistent with all lines

Pitch -0.0022530
Roll -0.0010315

comment:12 Changed 4 years ago by asm

Lidar Processing
Lidar classification started.

comment:13 Changed 4 years ago by asm

Lidar Processing
Lidar classification finished.

comment:14 Changed 4 years ago by asm

Lidar Processing
Elevation difference measured manually. Waiting for making delivery.

comment:15 Changed 4 years ago by asm

Lidar Processing
Script make_arsf_delivery has been updated. Delivery and readme file created.

Status: Done. Ready for DC.

comment:16 Changed 4 years ago by lah

Lidar DC

  • Needs a lidar quality report
  • Have removed empty Eagle & Hawk page from logsheet
  • converted ASCII files to CLRF terminated
  • pitch & roll look ok
  • classification mostly ok. Some missed points below ground on line 3, but probably not worth redoing
  • copied in sol file
  • Have moved extra intensity mosaic out of delivery and into processing folder as I don't think this is needed
  • demcompare masked mean 4.69
  • fw files look fine in waveviewer

Ready to go when lidar quality report is finalised and added.

comment:17 Changed 4 years ago by asm

Lidar DC

Dan has created the new quality report and I have added to the delivery.

Ready to go

comment:18 Changed 4 years ago by asm

Data Delivery
Mail sent to PI and waiting in order to let him select delivery method.

comment:19 Changed 4 years ago by asm

Lidar Data Delivery
PI selected FTP server, created and delivered. All clear for Lidar.

comment:20 Changed 4 years ago by asm

Hyperspectral Processing
Delivery created. It still needs the updtated data quality report but the project is ready for starting with Delivery Check.

P.S.: New data quality report has been added to the delivery

Last edited 4 years ago by asm (previous) (diff)

comment:21 Changed 4 years ago by gej

Hyperspectral Delivery Check

ReadMe:
apltran -inproj latlong WGS84 -igm f169b013b.igm -output
f169b023b_utm_wgs84N31.igm -outproj utm_wgs84N 31
should be:
apltran -inproj latlong WGS84 -igm f169b023b.igm -output
f169b023b_utm_wgs84N31.igm -outproj utm_wgs84N 31

Need to fix in config file aswell.

Everything else looks okay.
Have checked spectra and looks fine to me.

comment:22 Changed 4 years ago by asm

Hyperspectral Delivery Check

Have made the changes on the config file, re-created the readme and replaced on the delivery folder.
Problem found with SCT value on flightline 14. Will check.

Last edited 4 years ago by asm (previous) (diff)

comment:23 Changed 4 years ago by asm

Hyperspectral Delivery Check
Have created a new mapped file for flighltine 14. Updated screenshot, mosaic and Readme file.
Ready for resuming DC.

comment:24 Changed 4 years ago by gej

Hyperspectral Delivery Check

Check flightline 14, looks fine. Screenshots and mosaic look fine.
Readme has been updated and apltran command fixed.

Ready for delivery

comment:25 Changed 4 years ago by gej

Hyperspectral Data Delivered to PI on 20th October 2015 via FTP

comment:26 Changed 3 years ago by dac

Archiving

Started archiving

comment:27 Changed 3 years ago by gej

PI had problems with download data via FTP so has now been delivered via HDD to University of Balearic Islands, 18th November 2015

comment:28 Changed 3 years ago by dac

  • Resolution set to fixed
  • Status changed from new to closed

Archiving

Data now available via NEODC.

comment:29 Changed 3 years ago by dac

LiDAR

Started reprocessing with full range of scan angles.

comment:30 Changed 3 years ago by dac

LiDAR

Files ran through OK (discrete and full waveform) classifying now.

comment:31 Changed 3 years ago by dac

LiDAR

Classified data - creating delivery.

comment:32 Changed 3 years ago by dac

LiDAR

Created delivery - ready for check.

comment:33 Changed 3 years ago by stgo

Lidar delivery check
starting check

comment:34 Changed 3 years ago by stgo

Lidar delivery check

Delivery is fine. AGC striping in the screenshots is pretty bad, I know we already mention this in the readme but is it worth making a specific comment to explain how it effects the data visually? If not this is ready to go!

comment:35 Changed 3 years ago by dac

Added the following to the readme:

The recorded intensity exhibits more variation than normal for this flight due to the automatic gain control. This is visible as stripes along-track in the intensity image.

Will deliver via hard drive.

comment:36 Changed 3 years ago by dac

LiDAR

Reprocessed LiDAR data sent on hard drive (15/03/2016).

comment:37 Changed 3 years ago by dac

Archiving

Started uploading reprocessed LiDAR data to NEODC.

comment:38 Changed 3 years ago by dac

Archiving

Reprocessed LiDAR data uploaded to NEODC - will notify Wendy once 170 is reprocessed.

comment:39 Changed 3 years ago by dac

Archiving

Reprocessed LiDAR now available from NEODC.

Note: See TracTickets for help on using tickets.